OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet? A. B. C. D.

OpenStudy (anonymous):

A. 4x-3 B. 2x3 C. 2x2 -6x D. 2X3-6X2

OpenStudy (anonymous):

Volume of a cuboid= l * b * h Width = x Length = l = 2x Height = h = x - 3 Now substitute the values in the formula. \[2x \times x \times (x-3)\] Can you solve it now?

OpenStudy (anonymous):

4X-3

OpenStudy (anonymous):

No, \[2x ^{2} ( x - 3) \] Therefore, it's \[2x ^{3} - 6x ^{2}\]

OpenStudy (e.mccormick):

Tip: Volume is always a cubic equation, so it must have something to the third power. In a question like this, little facts light that can eliminate some of the choices and make it easier to get to the right answer.

OpenStudy (anonymous):

OH I get it now, thanks a lot

OpenStudy (anonymous):

@e.mccormick Indeed, if there are choices that is. :-)

OpenStudy (e.mccormick):

@Jas9420 Yep. And it can also help you see if you did a calculation right. if you multiply out a volume and don't get a cubic value, then you made a mistake somewhere!